what is the surface area of a piece of pipe that is open at both ends , has a diameter of 4 feet , and has a height of 12.5 feet ? and we must show our work
Answer by checkley77(12844)

C=2PIR is the circumferance.
C=2*3.14*4/2
C=6.28*2=12.56 is the ciecumferance.
SA=C*L
SA=l12.56*12.5=157 ft^2 is the surface area.